Kb for an unknown base is 3.7 x 10-9. What is the acid dissociation constant for the conjugate acid?
1
Expert's answer
2020-06-10T07:37:26-0400

"Ka" is the acid dissociation constant.


"Ka*Kb=Kw"


where kw is the water dissociation constant.

and "Kw" =1.0 * 10^-14


"Ka*(3.7*10^{-9})=1.0*10^{-14}"

"Ka=\dfrac{1.0*10^{-14}}{3.7*10^{-9}}"


"=2.7*10^{-6}"